OpenStudy (anonymous):

Given the function f(x) = x2 + 8x + 14 A Rewrite the function in vertex form by completing the square.

OpenStudy (anonymous):

complete the square you get (x+4)^2-2

OpenStudy (anonymous):

i'll show you a great vid

OpenStudy (anonymous):

how di d you get -2 i got + 30?

OpenStudy (anonymous):

\[f(x) = x^{2} + 8x + 14\]\[x^{2} + 8x + 14 = 0\]\[x^{2} + 8x = -14\]\[x^{2} + 8x + 16 = -14 + 16\]\[(x + 4)^{2} = 2\]\[0 = (x + 4)^{2} - 2\]\[f(x) = (x + 4)^{2} - 2\]

OpenStudy (anonymous):

Welcome to Open Study were everyone is here is smart lol hahah just kidding.. bt anyway welcome XD

OpenStudy (anonymous):

thank you so much!!

OpenStudy (anonymous):

np :)